Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.15 (given its target capital structure). Vandell has $11.41 million in debt that trades at par and pays an 7.3% interest rate. Vandell’s free cash flow (FCF0) is $1 million per year and is expected to grow at a constant rate of 6% a year. Vandell pays a 30% combined federal and state tax rate. The risk-free rate of interest is 3% and the market risk premium is 8%. Hastings’ first step is to estimate the current intrinsic value of Vandell. What are Vandell’s cost of equity and weighted average cost of capital? Round your answer to two decimal places. Do not round intermediate calculations. Cost of equity: % WACC: % What is Vandell's intrinsic value of operations? (Hint: Use the free cash flow corporate valuation model.) Round your answer to two decimal places. Do not round intermediate calculations. $ million What is the current intrinsic value of Vandell's stock? Round your answer to the nearest cent. Do not round intermediate calculations. $ /share
Answer #1

A)The calculation of Vandell’s cost of equity and weighted average cost of capital is shown below:

Cost of equity = Risk free rate + beta * market premium = 3% + 1.15 * 8% = 12.2%

After tax cost of debt = before tax cost of debt * (1 - tax rate) = 7.3% * (1-30%) = 5.11%

WACC = (Cost of equity + weight of equity ) + (cost of debt * weight of debt) = (70% * 12.2%) + (30% *5.11%)

= 10.07%

B) The calculation of Vandell's intrinsic value of operations is shown below:

value of operation = ((1 + constant growth rate)*FCF0)/(WACC - constant growth rate)

= ((1+6%)*$1,000,000)/(10.07% - 6%)

= $26,044,226.04

C)The calculation of current intrinsic value of Vandell's stock is shown below:

Value of debt = $11,410,000

Value of equity = $26,044,226.04 - $11,410,000 = $14,634,226

Number of shares = 1,000,000 shares

Intrinsic value per share = $14,634,226/1,000,000 shares = $14.63